# Operational Methods Applied to the Spherical Mean and X-Ray Transform

### v3

@@ Abstract
We use the framework of operational calculus to derive the operators associated with the spherical mean and related means of a function in $n$-dimensional space. These operators are then used to simplify the treatment and discussion of the means, i. e., to determine their power series expansion, to find the series solution to the corresponding inversion problem, to derive the associated differential equations that the means satisfy, or give meaning to fractional spherical means. Finally, we use the framework to derive the inversion formula of the X-ray transform in an operational manner.
@@ 1 Spherical Mean and Its Generalizations
The general solution of the wave equation in odd space dimensions is given by Kirchhoff’s formula and derived by using spherical means to reduce the higher dimensional wave equation to the wave equation in $\mathbb{R}$, and then using d’Alembert’s famous formula to solve the reduced system [11]. Spherical means also appear in the theory of harmonic functions as one of their defining properties, i. e. a function is harmonic if and only if it is its own spherical mean [18]. The spherical mean of a function $f$ in $n$ dimensions over a sphere of radius $r$ is defined by the integral [12]
@@
$\mathrm{e}^{t\hat{D}_{x}}f(x)=f(x+t).$ (2)
This expression can be easily generalized to higher dimensions by using the usual properties of the exponential function and the commutativity of the various derivative operators, giving
@@
As mentioned above, this relation can be used to find a compact (though somewhat unusual) expression for the spherical mean as an operator. We find by directly applying Eq. (3) to equation (1):
@@
The integral on the right-hand side is independent of $f(\vec{x})$ and can be calculated using standard procedures [8], yielding the operator expression
@@
where $I_{\alpha}(x)$ is the modified Bessel function of the first kind [1]. In fact, we can write this expression even more compact by using generalized hypergeometric functions [2],
@@
or more precisely, the confluent hypergeometric limit function, as
@@
This operator is a function of the Laplacian $\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}$ acting on the function $f(\vec{x})$ and contains all relevant information about the spherical mean of that function. To make sense of the operator, we interpret it in terms of the power series of $_{0}F_{1}(;\frac{n}{2};x)$ at $x=0$:
@@
This already gives us a way to efficiently approximate the spherical mean for small radii $r$. The same expression was found using standard methods and Taylor’s theorem in Ref. [20], confirming this result is correct. Before we continue, let us emphasize that the operational methods are not rigorous and therefore need to be verified by conventional techniques. However, as shown above, they provide valuable insights and “shortcuts” for finding analytical results that are otherwise more cumbersome to derive. With the help of the spherical mean operator, it is an easy task to write down the inverse operation of spherical averaging by simply diving by $_{0}F_{1}(;\frac{n}{2};\frac{r^{2}\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}{4})$ on both sides:
@@
Again, we interpret the operator by its power series and find the series solution to the inversion problem as
@@
Furthermore, successive spherical means of the same function can be found by applying the operator (7) multiple times to the function. Looking at the $m$-th iterated spherical mean, we find the power series expansion for small $r$:
@@
This formula also is valid for $m<0$ and therefore generalizes Eq. (9). By continuation, this expression also provides a way to compute and give meaning to fractional spherical means, such as the $\frac{1}{2}$-th spherical mean” of $f(\vec{x})$. In figure 1 we showcase different fractional spherical means in one dimension. From the operator form it is also clear that the successive spherical means of different radii $r_{1},r_{2}$ of a function commute, i. e.
@@
since the operators commute. Nonetheless, spherical means are not additive in the sense that $_{0}F_{1}\big{(};\frac{n}{2};\frac{r_{2}^{2}\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}{4}\big{)}\bar{f}(r_{1},\vec{x})\neq\bar{f}(r_{1}+r_{2},\vec{x})$. However, they admit an addition formula in terms of the Gaussian hypergeometric function $_{2}F_{1}$[5]
@@
Note that for $r_{1}=r_{2}=r$, two consecutive spherical means of equal radius, we get the rather simple hypergeometric expression [4]
@@
Figure 1: Demonstration of different spherical means in 1D. (a) Comparison of the spherical mean by the integral Eq. (1) (solid red line) with the spherical mean by the operator Eq. (7) (dashed red line) for the same radius of an arbitrary function $f(x)$. (b) Demonstration of various fractional spherical means. The solid blue line is the $\frac{1}{2}$-th spherical mean and the solid green line is the $\frac{1}{3}$-rd spherical mean. For comparison, the full spherical mean in red.
For the next paragraph, let us remember another property of harmonic functions. These functions, which are their own spherical mean, are known to satisfy Laplace’s equation:
@@
for appropriate boundary conditions on $\Omega$ [3]. While this is not true for the spherical mean of a function in general, it does satisfy a modified version of Laplace’s equation where we include additional source terms. To derive the corresponding differential equation, recall that the confluent hypergeometric limit functions $w\equiv{}_{0}F_{1}(;n;x)$ satisfy the following differential equation [2]:
@@
To bring this into the form of our operator, we perform the change of variables $x\to ax^{2}$ and write $s\equiv ax^{2}$, which gives us
@@
Applying the chain rule to the differential operators, we continue to find for the first derivative:
@@
Therefore, the differential equation (16) under the change of variables becomes
$\frac{\mathrm{d}^{2}w(ax^{2})}{\mathrm{d}{x}^{2}}+\frac{2n-1}{x}\frac{\mathrm{d}w(ax^{2})}{\mathrm{d}x}-4aw(ax^{2})=0,$ (20)
and by treating the Laplacian as a simple number, we can identify $a\to\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}/4$, which implies that the spherical mean operator “satisfies” the partial differential equation
@@
By applying this equation from the left to the function $f(\vec{x})$ and combining the operator and the function to the spherical mean, we obtain the well-known Euler-Poisson-Darboux equation [10, 23]:
@@
Now that we have seen the effectiveness and use of operational calculus and the operator associated with an integral transformation, let us consider other examples. As a matter of fact, the spherical mean can be seen as a special member of a broader class of averaging operations on a function within a spherical region of space:
@@
where $K(\vec{u})$ is the kernel or weight of the mean, which we assume to be radially symmetric, and $C_{K}$,
@@
is a normalization constant such that the average of the identity returns the identity. In this notation, the usual spherical mean is given by the choice of kernel $K(\vec{u})=\delta(\lVert\vec{u}\rVert-1)$. To find the associated operator of Eq. (23), recall the expression for the Fourier transform of radially symmetric functions [13]:
$\mathcal{F}\{f(\lVert\vec{x}\rVert)\}(k)=\frac{(2\pi)^{\frac{n}{2}}}{k^{\frac{n}{2}-1}}\int_{0}^{\infty}\mathrm{d}u\!\mathop{}J_{\frac{n}{2}-1}(ku)u^{\frac{n}{2}}f(u),$ (25)
using the shift operator (3) on our integral,
@@
the right-hand side is in essence the Fourier transform of $K(\lVert\vec{u}\rVert)\Theta(1-\lVert\vec{u}\rVert)$ after a “Wick rotation” in frequency space $k\to\mathrm{i}r$. Inserting the expression for the Fourier transform for the integral, we find
@@
from which we can further recognize the term $(r\lVert\vec{\nabla}_{\vec{x}}\rVert)^{-\frac{n}{2}+1}I_{\frac{n}{2}-1}(ru\lVert\vec{\nabla}_{\vec{x}}\rVert)$ from the ordinary spherical mean. Therefore, the integral could be equivalently written in terms of the confluent hypergeometric limit function:
$\bar{f}_{K}(r,\vec{x})=\frac{2\pi^{\frac{n}{2}}}{C_{K}\Gamma(\frac{n}{2})}\int_{0}^{1}\mathrm{d}u\!\mathop{}K(u)u^{n-1}{}_{0}F_{1}\Big{(};\frac{n}{2};\frac{(ru)^{2}\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}{4}\Big{)}f(\vec{x})$ (28)
and so it is easy to see that the generalized spherical mean for some kernel $K(u)$ is just a weighted integral of the spherical mean of $f(\vec{x})$:
@@
As a first simple example, consider the constant kernel $K(\vec{u})=1$; this gives the average of a function over the entire $n$-ball instead of just its surface. By equation (27) we find the following operator:
@@
Again, this operator can be used to find small $r$ approximations or the series solution to the inversion problem. Furthermore, since its form is almost identical to the spherical mean operator, we can immediately write down the corresponding partial differential equation the $n$-ball mean satisfies:
@@
A more interesting example is the general kernel $K(\vec{u})=(1-\lVert\vec{u}\rVert^{\beta})^{\alpha}$ with normalization constant $C_{K}=\frac{\Gamma(\alpha+1)\Gamma(\frac{n}{\beta})}{\beta\Gamma(\frac{n}{\beta}+\alpha+1)}S_{n-1}$. Let us consider two special cases. First, the bell-shaped $\beta=2$. Inserting the kernel into equation (27) we obtain the familiar looking operator
@@
which satisfies the differential equation
@@
Both expressions converge to the $n$-ball mean for $\alpha\to 0$, as expected. Note that using the property of hypergeometric function that $_{0}F_{1}(;a;x)=\frac{x}{(a+1)a}{}_{0}F_{1}(;a+2;x)+{}_{0}F_{1}(;a+1;x)$, we arrive at the peculiar relationship between the average over a sphere, $\bar{f}_{S}$, the average over a ball, $\bar{f}_{B}$, and the average over a shifted parabola, $\bar{f}_{K}$ ($\alpha=1$):
@@
The second case, $\beta=1$, is a generalization of the triangular weight and results in the operator
@@
Since the partial differential equation is more involved here, we will refrain from showing it. Again, these expressions converge to the $n$-ball mean for $\alpha\to 0$.
Before we move on to the next section, let us briefly consider two more examples. The first is the one-dimensional form of Eq. (30), which coincides with the moving average of a function over a window of width $r$:
@@
where by use of the two integral identities [6]
@@
we find for the inverse operation $f(x)=\frac{r\hat{D}_{x}}{\sinh(r\hat{D}_{x})}\bar{f}(r,x)$ that recovers the original function from the moving average:
@@
for suitable conditions on $\bar{f}(r,x)$.
The second example is related to the two-dimensional spherical mean,
@@
by “complexifying” the $y$-component of $\vec{u}$, yielding the integral
@@
However, applying the shift operator to the integrand shows that
@@
since $\int_{0}^{2\pi}\mathrm{d}\theta\!\mathop{}\mathrm{e}^{r\mathrm{e}^{\mathrm{i}\theta}}=2\pi$ for $r\neq 0$. So the “complex” 2D spherical mean of a function is the function itself; this is known as Gauss’s mean value theorem [22]. Obviously this is not a valid proof, since we have omitted the necessary conditions on $f$, but it is nevertheless a simple way of arriving at this conclusion.

### v4

@@ Abstract
We employ the framework of operational calculus to derive the operators associated with the spherical mean and a class of related averaging means of a function in $n$-dimensional space. Beginning with the classical definition of the spherical mean, we obtain a compact operator representation in terms of confluent hypergeometric functions of the Laplacian. This operator-based formulation provides a straightforward approach to the analysis of spherical means, allowing us to determine their power series expansions, construct series solutions to the corresponding inversion problems, derive the partial differential equations they satisfy, and give meaning to iterated and fractional spherical means. Finally, we apply the spherical mean operator to derive the inversion formula for the X-ray transform in an operational manner.
Keywords: Operational methods; Spherical mean; X-ray transform; Hypergeometric functions
@@ 1 Spherical Mean and Its Generalizations
The general solution of the wave equation in odd spatial dimensions is given by Kirchhoff’s formula, which is derive by employing spherical means to reduce the higher-dimensional wave equation to the wave equation in $\mathbb{R}$ and then applying d’Alembert’s celebrated formula to solve the resulting reduced system [11]. Spherical means also arise naturally in the theory of harmonic functions as one of their defining characteristics; specifically, a function is harmonic if and only if it coincides with its own spherical mean [18]. The spherical mean of a function $f$ in $n$ dimensions over a sphere of radius $r$ is defined by the integral [12]
@@
$\mathrm{e}^{t\frac{\partial}{\partial x}}f(x)=f(x+t).$ (2)
We can readily generalize this expression to higher dimensions by invoking the standard properties of the exponential function together with the commutativity of the individual derivative operators, yielding
@@
As mentioned above, we can use this relation to obtain a compact, albeit somewhat unconventional, operator expression for the spherical mean. By directly applying Eq. (3) to Eq. (1), we find
@@
The integral on the right-hand side is independent of the specific form of $f(\vec{x})$ and can be evaluated using standard techniques [8], which yields the operator representation
@@
where $I_{\alpha}(x)$ denotes the modified Bessel function of the first kind [1]. We can express this result in an even more compact form by using generalized hypergeometric functions [2],
@@
or, more specifically, the confluent hypergeometric limit function, which allows us to write
@@
This operator is a function of the Laplacian $\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}$ acting on $f(\vec{x})$ and therefore contains all relevant information about the spherical mean of the function. To give this operator a precise meaning, we interpret it through the power series expansion of $_{0}F_{1}(;\frac{n}{2};x)$ about $x=0$:
@@
This expansion immediately provides an efficient means of approximating the spherical mean for small radii $r$. Using standard techniques based on Taylor’s theorem, Ref. [20] obtained the same expression, thereby confirming the validity of this result. Before proceeding further, we emphasize that operational methods are not rigorous and must therefore be verified using conventional analytical techniques. Nevertheless, as demonstrated above, they offer valuable insights and practical “shortcuts” for deriving analytical results that would otherwise be considerably more involved. With the spherical mean operator at hand, we can readily write down the inverse operation of spherical averaging by dividing both sides by $_{0}F_{1}(;\frac{n}{2};\frac{r^{2}\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}{4})$:
@@
Interpreting this inverse operator via its power series expansion, we obtain the corresponding series solution to the inversion problem as
@@
Moreover, we can compute successive spherical means of the same function by applying the operator (7) repeatedly. Considering the $m$-th iterated spherical mean, we obtain the following power series expansion for small $r$:
@@
This formula remains valid for $m<0$ and thus generalizes Eq. (9). By analytic continuation, it also provides a natural framework for defining and computing fractional spherical means, such as the $\frac{1}{2}$-th spherical mean” of $f(\vec{x})$. In Fig. 1, we illustrate several fractional spherical means in one spatial dimension. From the operator formulation, it is also evident that successive spherical means with different radii $r_{1}$ and $r_{2}$ commute, that is,
@@
since the underlying operators commute. Nevertheless, spherical means are not additive in the sense that $_{0}F_{1}\big(;\frac{n}{2};\frac{r_{2}^{2}\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}{4}\big)\bar{f}(r_{1},\vec{x})\neq\bar{f}(r_{1}+r_{2},\vec{x})$. Instead, they satisfy an addition formula involving the Gaussian hypergeometric function $_{2}F_{1}$[5]
@@
In particular, for $r_{1}=r_{2}=r$, corresponding to two consecutive spherical means of equal radius, we obtain the comparatively simple hypergeometric representation [4]
@@
Figure 1: Demonstration of different spherical means in 1D. (a) Comparison of the spherical mean obtained from the integral in Eq. (1) (solid red line) with the spherical mean obtained from the operator in Eq. (7) (dashed red line) for the same radius applied to an arbitrary function $f(x)$. (b) Illustration of various fractional spherical means. The solid blue line represents the $\frac{1}{2}$-th spherical mean, and the solid green line represents the $\frac{1}{3}$-rd spherical mean. For reference, the full spherical mean is shown in red.
For the next paragraph, we recall another fundamental property of harmonic functions. These functions, which coincide with their own spherical mean, are known to satisfy Laplace’s equation,
@@
subject to appropriate boundary conditions on $\Omega$ [3]. Although this property does not generally hold for the spherical mean of an arbitrary function, the spherical mean does satisfy a modified version of Laplace’s equation that includes additional source terms. To derive the corresponding differential equation, we recall that the confluent hypergeometric limit functions $w\equiv{}_{0}F_{1}(;n;x)$ satisfy the ordinary differential equation [2]
@@
To bring this equation into a form compatible with our operator, we perform the change of variables $x\to ax^{2}$ and introduce the notation $s\equiv ax^{2}$, which yields
@@
Applying the chain rule to the differential operators, we obtain for the first derivative
@@
Consequently, under this change of variables, the differential equation (16) becomes
$\frac{\mathrm{d}^{2}w(ax^{2})}{\mathrm{d}{x}^{2}}+\frac{2n-1}{x}\frac{\mathrm{d}w(ax^{2})}{\mathrm{d}x}-4aw(ax^{2})=0.$ (20)
By formally treating the Laplacian as a scalar quantity, we can identify $a\to\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}/4$, which implies that the spherical mean operator formally “satisfies” the partial differential equation
@@
Applying this equation from the left to the function $f(\vec{x})$ and combining the operator with the function to recover the spherical mean, we obtain the well-known Euler–Poisson–Darboux equation [10, 23]:
@@
Having demonstrated the effectiveness and utility of operational calculus and the operator associated with an integral transformation, we now turn to additional examples. In fact, we can regard the spherical mean as a particular instance of a broader class of averaging operations applied to a function over a spherical region of space:
@@
where $K(\vec{u})$ denotes the kernel, or weight, of the mean, which we assume to be radially symmetric, and $C_{K}$,
@@
is a normalization constant chosen such that averaging the identity function returns the identity. Within this framework, the standard spherical mean corresponds to the choice of kernel $K(\vec{u})=\delta(\lVert\vec{u}\rVert-1)$. To determine the operator associated with Eq. (23), we recall the expression for the Fourier transform of radially symmetric functions [13]:
$\mathcal{F}\{f(\lVert\vec{x}\rVert)\}(k)=\frac{(2\pi)^{\frac{n}{2}}}{k^{\frac{n}{2}-1}}\int_{0}^{\infty}\mathrm{d}u\!\mathop{}J_{\frac{n}{2}-1}(ku)u^{\frac{n}{2}}f(u).$ (25)
Using the shift operator (3) in the integral representation,
@@
we observe that the right-hand side is essentially the Fourier transform of $K(\lVert\vec{u}\rVert)H(1-\lVert\vec{u}\rVert)$ after a “Wick rotation” in frequency space $k\to\mathrm{i}r$. Substituting the Fourier transform expression into the integral, we obtain
@@
from which we recognize the factor $(r\lVert\vec{\nabla}_{\vec{x}}\rVert)^{-\frac{n}{2}+1}I_{\frac{n}{2}-1}(ru\lVert\vec{\nabla}_{\vec{x}}\rVert)$ familiar from the ordinary spherical mean. Consequently, we can equivalently express the integral in terms of the confluent hypergeometric limit function:
$\bar{f}_{K}(r,\vec{x})=\frac{2\pi^{\frac{n}{2}}}{C_{K}\Gamma(\frac{n}{2})}\int_{0}^{1}\mathrm{d}u\!\mathop{}K(u)u^{n-1}{}_{0}F_{1}\Big(;\frac{n}{2};\frac{(ru)^{2}\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}{4}\Big)f(\vec{x}),$ (28)
which makes it evident that the generalized spherical mean associated with a kernel $K(u)$ is simply a weighted integral of the spherical mean of $f(\vec{x})$:
@@
As a first simple example, we consider the constant kernel $K(\vec{u})=1$, which yields the average of a function over the entire $n$-ball rather than only over its surface. From Eq. (27), we directly obtain the corresponding operator:
@@
As before, we can employ this operator to derive small-$r$ approximations or to construct the series solution of the associated inversion problem. Moreover, because its structure closely resembles that of the spherical mean operator, we can immediately write down the partial differential equation satisfied by the $n$-ball mean:
@@
A more interesting example arises from the general kernel $K(\vec{u})=(1-\lVert\vec{u}\rVert^{\beta})^{\alpha}$ with normalization constant $C_{K}=\frac{\Gamma(\alpha+1)\Gamma(\frac{n}{\beta})}{\beta\Gamma(\frac{n}{\beta}+\alpha+1)}S_{n-1}$. We now examine two special cases. First, we consider the bell-shaped case $\beta=2$. Substituting this kernel into Eq. (27), we obtain the familiar operator
@@
which satisfies the partial differential equation
@@
Both expressions converge to the $n$-ball mean in the limit $\alpha\to 0$, as expected. Furthermore, by using the hypergeometric identity $_{0}F_{1}(;a;x)=\frac{x}{(a+1)a}{}_{0}F_{1}(;a+2;x)+{}_{0}F_{1}(;a+1;x)$, we obtain the following notable relationship between the spherical average $\bar{f}_{S}$, the ball average $\bar{f}_{B}$, and the average over a shifted parabola $\bar{f}_{K}$ (with $\alpha=1$):
@@
The second case, $\beta=1$, generalizes the triangular weight and leads to the operator
@@
The associated partial differential equation is considerably more involved and given by
$\frac{2r}{n+1}\frac{\mathrm{d}^{3}\bar{f}_{K}}{\mathrm{d}{r}^{3}}+\frac{2(n+\alpha)}{n+1}\frac{\mathrm{d}^{2}\bar{f}_{K}}{\mathrm{d}{r}^{2}}+\left[\frac{\alpha^{2}+n(n-1)+(2n-1)\alpha}{(n+1)r}+\frac{r\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}{(n+1)}\right]\frac{\mathrm{d}\bar{f}_{K}}{\mathrm{d}r}=\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}\bar{f}_{K}.$ (36)
Once again, these expressions reduce to the $n$-ball mean in the limit $\alpha\to 0$.
Before proceeding to the next section, we briefly discuss two additional examples. The first concerns the one-dimensional version of Eq. (30), which coincides with the moving average of a function over a window of width $r$:
@@
where, by employing the two integral identities [6]
@@
we obtain the inverse operation $f(x)=\frac{r\hat{D}_{x}}{\sinh(r\hat{D}_{x})}\bar{f}(r,x)$, which reconstructs the original function from its moving average:
@@
under suitable regularity conditions on $\bar{f}(r,x)$.
The second example relates to the two-dimensional spherical mean,
@@
by “complexifying” the $y$-component of $\vec{u}$, which leads to the integral
@@
Applying the shift operator to the integrand, however, shows that
@@
since $\int_{0}^{2\pi}\mathrm{d}\theta\!\mathop{}\mathrm{e}^{r\mathrm{e}^{\mathrm{i}\theta}}=2\pi$ for $r\neq 0$. Consequently, the “complex” two-dimensional spherical mean of a function coincides with the function itself, a result known as Gauss’s mean value theorem [22]. Although this argument does not constitute a rigorous proof because it omits the necessary conditions on $f$, it nevertheless provides an intuitive and concise route to this classical conclusion.